OpenStudy (anonymous):

Given f(x) = 2 x^2 + 4 x + 5 , determine the point on the curve where the tangent line to the curve is horizontal. (The answer must be a POINT on the curve.) The point on the curve where the tangent line is horizontal is (). (Separate the coordinate values with a comma.)

OpenStudy (anonymous):

Differentiate, set the derivative to zero, solve for x. Substitute that value into f to get the appropriate y value. Show answer as (x,y).

OpenStudy (anonymous):

thanks, this was helpful

OpenStudy (anonymous):

(x,Y) = (-1,3)

OpenStudy (anonymous):

\[f(x) = 2 x^2 + 4 x + 5\]\[f'(x)=4x+4=0\rightarrow x=-1\rightarrow f(-1)=3\]So the solution is (x.y)=(-1,3)

OpenStudy (anonymous):

I apparently don't keep up on typing skills.....

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

By the way, you could use algebra to find the x coordinate of the vertex (x=-b/2a) to get the same result with less sophistication.

OpenStudy (anonymous):

cool, good to know for the future :)

OpenStudy (anonymous):

Of course, that only works on quadratics; the calculus method works in general.

OpenStudy (anonymous):

ok